Project

General

Profile

NDOS CommissionMtng 20120723

Attending: Martin, Peter, Athans, Jessica, Kurt, Jan, Susan Dominick, Nick,

UPDATES:
1.  Detector Update
Ran over the weekend from the daq cluster with vnc sessions on master.

2.  DataCheck Update
Fri-sat (Some loss of data)
saturday-sunday (smooth)

3 problem FEBs
2-1 21 (Shuts off
2-3 11 (16/19)

Previous week: 7/16 - 7/22
Monday to Tuesday
13977 18 subruns
FEB2-1:21 shutoff in subrun 0
FEB3-2:12 shutoff in subrun 10

Tuesday to Wednesday
13978 23 subruns
FEB2-1:21 shutoff in subrun 0
FEB2-3:39 shutoff in subrun 16
FEB1-1:38 shutoff in subrun 17

Wednesday to Thursday
13980 24 subruns
FEB2-1:21 shutoff in subrun 0
FEB3-1:0 shutoff in subrun 14

13981 2 subruns
Nothing to report

Thursday to Friday
13983 19 subruns
FEB3-2:23 shutoff in subrun 9
FEB1-1:38 shutoff in subrun 14
FEB2-3:11 shutoff in subrun 16
FEB2-3:32 shutoff in subrun 16

Friday to Saturday
13984 24 subruns
FEB2-1:21 shutoff in subrun 0
FEB3-2:23 shutoff in subrun 0
FEB3-2:48 shutoff in subrun 1
FEB3-2:51 shutoff in subrun 1

Saturday to Sunday
13985 24 subruns
Nothing to report

Sunday to Monday
13896 19 subruns
FEB2-3:11 shutoff in subrun 13

Problem FEBs
FEB2-1:21
FEB2-3:11
FEB3-2:23

WM APD Histories offline now (working on automated reprocessing)

4.  Installation Plans for the week
Lost 8 silicone APDs since lat week. Will try running them warm.
Caltech still working to get broken thermistors repaired

5.  DAQ Update and Plans for the week
Crookston didn't show up last week (Expected this week)
Kurt - Partitioning tests
Update DCM with new DCMApp to allow status readback with all pixels disabled
Mat Asked: Are we going to continue running the VNC session? A: Yes, except External webpages will run locally on nova-daq-04
Reny Scott might be able to help with VNC smoothness.

6.  DCS Update and Plans for the week
Monitors. Athans passed on budget code to Andrew. Monitors are being ordered through Vince.

7. Block Monitoring: Synoptic pages are being worked on.